# 03F40 (primary), 03D20

### v2

@@ 8 Finite Base Hierarchies
Note, however, that $\mathcal{D}$ in the above depends on $m$, so one may still ask if there is a single finite dynamical hierarchy which yields independence. Indeed, such a dynamical hierarchy can be constructed by diagonalization.
Theorem 2.17+
There is a finite, computable dynamical hierarchy $\mathcal{D}$ such that $\sf KP$ does not prove that if $m\in\mathbb{N}$, then there is $i\in\mathbb{N}$ such that $\mathbb{G}^{\mathcal{D}}_{i}(m)=0$.
Proof 8.2 (sketch).
@@
9 Concluding remarks
We have exhibited a Goodstein process based only on exponential notation, yet independent of very strong theories. This opens a new pathway to obtaining powerful extensions of Goodstein’s classic result. A somewhat less powerful Goodstein principle is obtained by using a version of the Ackermann function instead of the exponential [1]; it would be interesting to see if the combination of the two approaches leads to substantially more powerful principles, thus avoiding the use of Hardy functions in the search for even longer Goodstein processes.
Another interesting question we have not addressed is whether our processes retain the maximality condition of the Goodstein’s original, where the normal forms employed lead to the longest possible termination time among all processes using the same bases [10]. Here one may ask what ‘all processes’ should be considered for a maximality principle: should we consider variations of the dynamical hierarchy? Or, simply variations of notation within a fixed dynamical hierarchy?
Finally, we note a peculiarity of our proofs of termination and independence. We have used a different ordinal interpretation for each, but normally, one employs a single ‘optimal’ ordinal interpretation. We believe that it is possible to give such an interpretation, but computations become much more cumbersome. In particular, independence would use the fundamental sequences for $\vartheta$ [4], which are much less straightforward than those for $\psi$. Overall, we felt that the simplifications afforded by using two interpretations outweighed the cost of introducing a second collapsing function. While we do believe that a single interpretation could be used in our case, a dual-interpretation approach may prove indispensable in devising Goodstein processes independent of increasingly strong subsystems of second order arithmetic, perhaps one day including second order arithmetic itself.

### v4

@@ 8. Finite Base Hierarchies
Note, however, that $\mathcal{D}^{\prime}$ depends on $m$, so one may still ask if there is a single finite dynamical hierarchy which yields independence. Indeed, such a dynamical hierarchy can be constructed by diagonalization.
Theorem 2.
There is a finite, computable dynamical hierarchy $\mathcal{D}^{\star}$ such that $\sf KP$ does not prove that if $m\in\mathbb{N}$, then there is $i\in\mathbb{N}$ such that $\mathbb{G}^{\mathcal{D}}_{i}(m)=0$.
Proof sketch.
@@
9. Concluding remarks
We have exhibited a Goodstein process based only on elementary functions, yet independent of very strong theories. This opens a new pathway to obtaining powerful extensions of Goodstein’s classic result. A Goodstein principle of $\Gamma_{0}$-strength (much weaker than $\vartheta(\varepsilon_{{\Omega}+1})$) is obtained by using a version of the Ackermann function instead of the exponential, but with singleton base hierarchies [1]; it would be interesting to see if the combination of the two approaches leads to substantially more powerful principles, thus avoiding the use of Hardy functions in the search for ever more powerful Goodstein processes.
A second direction for future work lies in optimizing of our independence result. The primary source of complexity of fractal Goodstein processes lies in the base hierarchies involved; we have already established that only finitely many bases must be added at each step, but a simpler process may be obtained by adding bases more sparingly. Indeed, we added new bases for each $n$ that is a multiple of its own base, but as observed in Remark 7.3, this does not necessarily lead to $U_{B}(n)$ having uncountable cofinality. We could instead consider a dynamical base hierarchy which only adds new bases when $\tau(U_{B}(n))={\Omega}$, which would lead to independence with base hierarchies smaller than the greedy ones used. We have only resisted this approach due to it being less number-theoretically natural, but our proofs readily apply to such a variation. However, just how sparse we can make the base hierarchies while maintaining independence remains an interesting open question, possibly leading to phase transition results [15].
Finally, there is the issue of optimizing the ordinal assignment(s). Our goal was to keep ordinal calculations as simple as possible, which are already non-trivial with our approach. As we have mentioned, a tight ordinal interpretation would lead to more cumbersome computations. It would moreover not improve on the independence results for $\sf KP$, as these are not sensitive to the difference between $\vartheta$ and $\psi$. However, more fine-grained independence results could be obtained by a precise ordinal interpretation, such as independence for weaker theories with respect to restricted versions of our fractal Goodstein process, so a more refined ordinal analysis may be a viable direction for future work. For now, we find that the novelty of a two-interpretation approach has greater added value, as we predict that it will become indispensable as Goodstein principles are pushed farther beyond the limits of predicativity.
Acknowledgements
We thank Milan Morreel for many insightful comments and suggestions which have helped improve the article.